Disprove that the commutative property holds on the matrix multiplication.


1
Expert's answer
2021-10-12T10:23:44-0400

For example:

"A=\begin{pmatrix}
 3 & 4 \\
 1 & 2
\end{pmatrix}"


"B=\begin{pmatrix}
 6 & 2 \\
 3 & 2
\end{pmatrix}"


"AB=\begin{pmatrix}
 18+12 &6+8 \\
 6+6 & 2+4
\end{pmatrix}=\begin{pmatrix}
 30 &14 \\
 12 & 6
\end{pmatrix}"


"BA=\begin{pmatrix}
 18+2 &24+4 \\
 9+2 & 12+4
\end{pmatrix}=\begin{pmatrix}
 20 &28 \\
 11 & 16
\end{pmatrix}"


"AB\neq BA"